# Robust Spacecraft Low-Thrust Trajectory Design: A Chance-Constrained Covariance-Steering Approach

### v1

@@ IV-A Planar low-thrust Earth-to-Mars problem
We highlight that the effect of explicitly modeling mass dynamics, as a stochastic decision variable, was rigorously evaluated by comparing two formulations of the chance-constraint covariance-steering problem. In the first scenario, the mass equation was incorporated as a fifth state, and its uncertainty was propagated through the full closed-loop covariance dynamics. In the second scenario, mass evolution was considered deterministic and excluded from the set of decision variables, effectively assuming that the thrust-to-acceleration mapping remained constant throughout the mission. Despite producing comparable nominal trajectories, the results reveal notable discrepancies in the control effort and state uncertainty, particularly during the thrusting phases. Figure 7 compares the closed-loop feedforward control profiles for the two scenarios. The results show that including mass uncertainty leads to noticeably higher thrust magnitudes during the main thrust arcs, with a peak increase of approximately $6\%$ relative to the case without mass modeling.
@@ IV-B Three-dimensional (3D) Earth-to-Mars problem
The spacecraft dynamics are modeled in a Sun-centered Cartesian inertial frame, with the full state vector defined as $\bm{x}_{t}=[\,\bm{r}_{t}^{\top},\,\bm{v}_{t}^{\top},\,m_{t}\,]^{\top}\in\mathbb{R}^{7}$, where $\bm{r}_{t}=[r_{x},r_{y},r_{z}]^{\top}\in\mathbb{R}^{3}$, $\bm{v}_{t}=[v_{x},v_{y},v_{z}]^{\top}\in\mathbb{R}^{3}$ are the position and velocity vectors, respectively, and $m_{t}\in\mathbb{R}_{+}$ is the spacecraft mass. The dynamics are given by
@@
$d\begin{bmatrix}\bm{r}_{t}\\[3.0pt] \bm{v}_{t}\\[3.0pt] m_{t}\end{bmatrix}=\bm{f}(\bm{x}_{t},\bm{u}_{t},t)~dt+\begin{bmatrix}\bm{0}_{3\times 3}\\[6.0pt] \gamma I_{3}\\[6.0pt] 0\end{bmatrix}d\bm{w}_{t},$ (35)
where $d\bm{w}_{t}\in\mathbb{R}^{3}$ is a Wiener process and $\gamma$ is the constant acceleration disturbance intensity.
The continuous-time dynamics are linearized and discretized with $N=60$. The initial nominal trajectory is generated by solving a deterministic minimum-fuel problem using CasADi solver.
@@
The final state is constrained to $\sigma_{r_{f}}=3.16\times 10^{5}$ km and $\sigma_{v_{f}}=0.1$ km/s, with no explicit constraint on the final mass uncertainty, yielding
@@
Please note that a large, unpenalized variance is permitted for the final mass, reflecting that the control objective does not involve regulating the final mass uncertainty. The boundary conditions and physical parameters are summarized in Table II.
@@
The problem was scaled with a factor of $d=100$ for numerical conditioning and solved. The SCP termination criteria were set to $\varepsilon_{x}=5.0\times 10^{-4}$ and $\varepsilon_{\zeta}=1.0\times 10^{-6}$. The solver converged to a feasible solution in 12 iterations, demonstrating the method’s robustness and scalability to a problem with 7 states, 3 control inputs and with $N=60$ grid points.
@@
Unlike the position and velocity states, the initial mass is assumed to be known precisely ($\sigma_{m_{i}}=0$), reflecting accurate pre-launch propellant loading measurements. However, mass becomes a stochastic variable over time because the control input $\bm{u}_{t}$, which has an associated execution uncertainty, directly drives the mass rate equation $dm_{t}/dt\propto-\|\bm{u}_{t}\|$. This propagated mass uncertainty then couples back into the translational dynamics. The state transition matrix (STM) includes the sensitivity of acceleration to mass variations, $\partial(\bm{u}_{t}/m_{t})/\partial m_{t}=-\bm{u}_{t}/m_{t}^{2}$, creating a feedback loop where uncertainty in mass induces uncertainty in the achieved acceleration for a given thrust command.
@@
Figure 8: 3D view of the deterministic, open-loop trajectory for the Earth-to-Mars problem.
@@
The optimized controller successfully navigates the uncertain dynamics to meet the stringent terminal constraints. Figure 8 shows the resulting 3D nominal trajectory, illustrating the spacecraft’s path from Earth to Mars in the high-fidelity dynamical model. A Monte Carlo simulation with 1000 samples validates the controller’s performance and the accuracy of the covariance propagation. Figure 9 shows the $x$-$y$ projection of the closed-loop solution, where the position and velocity covariance ellipses are scaled by a factor of 10 for visual clarity. The close alignment of the Monte Carlo samples with the predicted $95\%$-confidence ellipses confirms the fidelity of the linearized covariance dynamics within the SCP loop. The control profile, shown in Figure 10, consists of three primary thrust arcs. The nominal thrust magnitude shows a modulated profile, deviating from a pure bang-bang structure to maintain the required robustness margins. The control covariance ellipsoids remain within the thrust upper limit, confirming chance constraint satisfaction.
(a) Closed-loop position covariances (10x).
(b) Closed-loop velocity covariances (10x).
@@
Figure 9: Monte Carlo validation (1000 samples) within the $95\%$-confidence ellipses.
Figure 10: Control thrust magnitude (top) and components (bottom) with associated $3\sigma$ uncertainty ellipsoids, demonstrating bounded stochastic control actions.
@@
One of our main analyses involves a direct comparison between the full stochastic-mass model and a simplified model where mass is treated deterministically. Figure 11 shows the time history of the mass and its standard deviation. The mean mass decreases by approximately $25\%$ over the course of the mission, reaching a final value of approximately $m_{f}=3775$ kg upon arrival at Mars. The mass standard deviation grows from zero to a final value of over 60 kg, driven entirely by the cumulative uncertainty in thrust execution.
@@
Figure 11: Time history of spacecraft mass and its standard deviation. The mass uncertainty grows from an initial value of zero due to the stochastic nature of the control inputs.
@@
This mass uncertainty has a notable impact on the translational states. Figure 12 shows the velocity standard deviation along the first principal component. The model with mass uncertainty (black) exhibits a peak velocity dispersion up to $17.73\%$ higher than the deterministic-mass model (red). This inflation occurs because mass uncertainty directly translates into acceleration uncertainty via the $1/m_{t}$ term in the dynamics. The effect on the overall position dispersion is summarized in Figure 13, which plots the trace of the position covariance matrix. Inclusion of mass uncertainty results in a $35.21\%$ higher peak position dispersion during the mid-course phase. This notable increase highlights that neglecting mass uncertainty leads to a significant underestimation of the required control authority and the true mission risk.
@@
Figure 12: Velocity standard deviation comparison. The model with mass uncertainty (black) shows significantly higher dispersion, especially during high-thrust phases.
@@
Figure 13: Trace of position covariance. The formulation with mass uncertainty (black) results in up to $35.21\%$ larger dispersion than the deterministic-mass model (red), highlighting the critical need to model this effect.
@@
Finally, Figure 14 compares the nominal control profiles. The controller for the full stochastic-mass model (black) commands up to $6\%$ higher thrust during the main arcs compared to the deterministic-mass controller (red). This additional control effort is necessary to counteract the larger predicted dispersions and actively shape the covariance, confirming that robustness to mass uncertainty carries a tangible fuel and performance cost.
@@
Figure 14: Comparison of nominal control thrust magnitudes. The controller accounting for mass uncertainty (black) requires higher thrust levels to achieve the same terminal covariance bounds as the controller that neglects it (red).
@@
This 3D case study conclusively demonstrates two key findings. First, the covariance-variable formulation is highly effective for robust trajectory optimization in complex, high-dimensional astrodynamics problems. Second, and more critically, the propagation of mass uncertainty has a substantial, quantifiable impact on both state dispersion and control effort. Neglecting this coupling, as is common in previously simplified models in the literature, leads to an underestimation of mission risk and an underpowered controller. The proposed framework successfully manages this coupled uncertainty, providing a robust and high-fidelity solution for generating robust interplanetary guidance algorithms.
@@ V Conclusions
This work presents a sequential convex programming framework for solving the chance-constraint covariance steering problem for low-thrust interplanetary transfers. The method incorporates a piecewise affine state-feedback policy and explicitly models spacecraft mass dynamics within the stochastic formulation. By coupling the control input and disturbance intensity through the time-varying spacecraft mass, the framework captures the evolution of uncertainty more realistically than conventional approaches.
Numerical results successfully demonstrate the regulation of both the mean trajectory and the state covariance, producing closed-loop solutions that satisfy probabilistic terminal constraints and remain consistent with Monte Carlo dispersion analysis. In particular, the inclusion of mass as a stochastic state reveals critical differences in both control effort and uncertainty evolution that are overlooked when mass is treated deterministically. Future studies will apply the computationally efficient proposed method to more challenging trajectory optimization problems [13].

### v2

@@ IV-A Planar (2D) low-thrust Earth-to-Mars problem
We highlight that the effect of explicitly modeling mass dynamics, as a stochastic decision variable, was rigorously evaluated by comparing two formulations of the chance-constraint covariance-steering problem. In the first scenario, the mass equation was incorporated as a fifth state, and its change was propagated through the full closed-loop covariance dynamics. In the second scenario, mass evolution was considered deterministic and excluded from the set of decision variables, effectively assuming that the thrust-to-acceleration mapping remained constant throughout the mission. Despite producing comparable nominal trajectories, the results reveal notable discrepancies in the control effort and state uncertainty, particularly during the thrusting phases. Figure 7 compares the feedforward control profiles ($\|\bm{F}\|$) for the two scenarios. The results indicate that including mass variations leads to noticeably higher thrust magnitudes during the main thrust arcs, with a peak increase of approximately $6\%$ relative to the case without mass modeling. Note that disturbance becomes stronger in the scenario that mass is decreased over the mission time.
@@ IV-B Three-dimensional (3D) Earth-to-Mars problem
The spacecraft dynamics are modeled in a Sun-centered Cartesian inertial frame, with the full state vector defined as $\bm{x}_{t}=[\,\bm{r}_{t}^{\top},\,\bm{v}_{t}^{\top},\,m_{t}\,]^{\top}\in\mathbb{R}^{7}$, where $\bm{r}_{t}=[r_{x},r_{y},r_{z}]^{\top}\in\mathbb{R}^{3}$, $\bm{v}_{t}=[v_{x},v_{y},v_{z}]^{\top}\in\mathbb{R}^{3}$ are the position and velocity vectors, respectively, and $m_{t}\in\mathbb{R}^{+}$ is the spacecraft mass. The dynamics are given by
@@
$d\begin{bmatrix}\bm{r}_{t}\\[3.0pt] \bm{v}_{t}\\[3.0pt] m_{t}\end{bmatrix}=\bm{f}(\bm{x}_{t},\bm{u}_{t},t)~dt+\begin{bmatrix}\bm{0}_{3\times 3}\\[6.0pt] \frac{\gamma}{m_{t}}I_{3}\\[6.0pt] 0\end{bmatrix}d\bm{w}_{t},$ (35)
where $d\bm{w}_{t}\in\mathbb{R}^{3}$ is a Wiener process and $\gamma$ is the constant acceleration disturbance intensity. The continuous-time dynamics are linearized and discretized with $N=60$. The initial nominal trajectory is generated by solving a deterministic minimum-fuel problem using CasADi solver.
@@
The final state is constrained to $\sigma_{r_{f}}=3.16\times 10^{5}$ km and $\sigma_{v_{f}}=0.1$ km/s, with a large value for the variance of the final mass, yielding
@@
Please note that a large, unpenalized variance is permitted for the final mass, $\sigma_{m_{f}}\leq\sqrt{5000}\approx 70.71$ kg, reflecting that the control objective does not involve regulating the dispersion in the final mass. The boundary conditions and physical parameters are summarized in Table II.
@@
The SCP termination criteria were set to $\varepsilon_{x}=5.0\times 10^{-4}$ and $\varepsilon_{\zeta}=1.0\times 10^{-6}$. The solver converged to a feasible solution in 12 iterations, demonstrating the method’s robustness and scalability to a problem with 7 states, 3 control inputs, and with $N=60$ grid points.
@@
Unlike the position and velocity states, the initial mass is assumed to be known precisely ($\sigma_{m_{i}}=0$), reflecting accurate pre-launch propellant loading measurements. However, mass becomes a stochastic variable over time because the control input $\bm{u}_{t}$, which has an associated execution uncertainty, directly drives the mass rate equation $dm_{t}/dt\propto-\|\bm{u}_{t}\|$. This propagated mass uncertainty then couples back into the translational dynamics. The state transition matrix (STM) includes the sensitivity of acceleration to mass variations, $\partial(\bm{u}_{t}/m_{t})/\partial m_{t}=-\bm{u}_{t}/m_{t}^{2}$, creating a feedback loop where uncertainty in mass induces uncertainty in the achieved acceleration for a given thrust command and also affects the strength of the disturbance.
@@
Figure 8: 3D view of the deterministic CasADi (black) and closed-loop (blue) trajectories for the Earth-to-Mars problem.
The optimized controller successfully navigates the uncertain dynamics to meet the stringent terminal constraints. Figure 8 shows the resulting 3D nominal trajectories, illustrating the spacecraft’s path from Earth to Mars in the more realistic 3D dynamical model. A Monte Carlo simulation with 1000 samples validates the controller’s performance and the accuracy of the covariance propagation.
@@
Figure 9 shows the $x$-$y$ projection of the closed-loop solution, where the position and velocity covariance ellipses are scaled by a factor of 10 for visual clarity. The close alignment of the Monte Carlo samples with the predicted $95\%$-confidence ellipses confirms the fidelity of the linearized covariance dynamics within the SCP loop. The control profile, shown in Figure 10, consists of three primary thrust arcs. Similar to the 2D case, the nominal thrust magnitude shows a modulated profile, deviating from a pure bang-bang structure to maintain the required robustness margins. The control covariance ellipsoids remain within the thrust upper limit, confirming chance constraint satisfaction.
(a) Closed-loop trajectory and position covariances (10x).
(b) Closed-loop velocity trajectory and covariances (10x).
@@
Figure 9: 3D case: Monte Carlo validation (1000 samples) within the $95\%$-confidence ellipses.
Figure 10: 3D case: magnitude of the Monte Carlo control samples (gray) and feed-forward nominal control (black).
@@
One of our main analyses involves a direct comparison between the full stochastic-mass model and a simplified model where mass is treated deterministically by using only the feed-forward control. Figure 11 shows the time history of the mass and its standard deviation, which is not necessarily a monotonically increasing value. In addition, the largest standard deviation occurs, in this case, at some point along the trajectory. If we enforce any constraint on the mass standard deviation at the final time to be greater than 57.7 kg ($\sigma_{m_{f}}\geq 57.7$ kg), this constraint is not going to become active, and the solution is more constrained by the final position and velocity variances. The mean mass decreases to a final value of approximately $m_{f}=3686.48$ kg upon arrival at Mars. As expected, the final mass is less than its 2D counterpart. The mass standard deviation grows from zero to a final value of nearly 57.7 kg, driven entirely by the cumulative uncertainty in thrust execution.
@@
Figure 11: 3D case: mass and its standard deviation vs. time. The standard deviation changes from an initial value of zero.
@@
This mass inclusion and its associated uncertainty has a notable impact on the translational states. Figure 12 shows the velocity standard deviation along the first principal component. The model with mass uncertainty (black) exhibits a peak velocity dispersion up to $34.84\%$ higher than the deterministic-mass model (red). This inflation occurs because mass uncertainty directly translates into acceleration uncertainty via the $1/m_{t}$ term in the dynamics. The effect on the overall position dispersion is summarized in Figure 13, which plots the trace of the position covariance matrix. Inclusion of mass uncertainty results in a $61.03\%$ higher peak position dispersion during the mid-course phase. This notable increase highlights that neglecting mass uncertainty leads to a significant underestimation of the required control authority and the true mission risk quantified in terms of the peak position and velocity variances.
@@
Figure 12: 3D case: Comparison of velocity standard deviations. The model with mass uncertainty (black) shows significantly higher dispersion, especially during the middle thrust arc.
@@
Figure 13: 3D case: trace of the position covariance. The formulation with mass uncertainty (black) results in up to $61.03\%$ larger dispersion than the deterministic-mass model (red).
@@
Figure 14 compares the nominal control profiles. The controller for the full stochastic-mass model (black) commands up to $6\%$ higher thrust during the main arcs compared to the deterministic-mass controller (red). This additional control effort is necessary to counteract the larger predicted dispersions and actively shape the covariance, confirming that robustness to mass uncertainty carries a tangible fuel and performance cost. Also, the middle thrust arc leverages the Oberth effect [33, 34] (i.e., it utilizes the gravitational well of the Sun to maximize the rate of change of its energy). Thus, it has a notable influence on the resulting position and velocity and their evolution along the trajectory.
@@
Figure 14: 3D case: comparison of the feed-forward control, $\|\bm{F}\|$. The controller accounting for mass uncertainty (black) requires higher thrust to achieve the same terminal covariance bounds compared to the controller that neglects it (red).
To further illustrate the utility of the proposed formulation, we solve the same 3D problem, but with
$P_{f}=\mathrm{diag}(10^{11},\ 10^{11},\ 10^{11},\ 10^{-2},\$
$10^{-2},\ 10^{-2},\ 1600)\quad\text{(km}^{2},\text{(km/s)}^{2},\text{kg}^{2}),$
which corresponds to $\sigma_{m_{f}}\leq 40$ kg (i.e., a maximum 40 kg standard deviation in final mass) at the end of the trajectory. Figure 15 shows the spacecraft mass and its standard deviation vs. time. An interesting observation is that while the standard deviation has decreased compared to the previous case, the final mass (3676.43) has decreased by 10 kg. More propellant is consumed to achieve the reduction of the dispersion in mass at the final time. We highlight that this result is similar to those obtained by solving optimal control problems using a desensitized optimal control formulation [24], which is another powerful method for generating robust optimal solutions [35, 36]. Figure 16 also compares the nominal thrust magnitude profiles for the previous case and when we enforce $\sigma_{m_{f}}\leq 40$ kg. The difference in the two profiles leads to the 10 kg more propellant consumption.
@@
The considered cases conclusively demonstrate two key findings. First, the covariance-variable formulation is highly effective for robust trajectory optimization in complex, high-dimensional astrodynamics problems. Second, and more critically, the propagation of mass uncertainty has a substantial, quantifiable impact on both state dispersion and control effort. Neglecting this coupling, as is common in previous models in the literature, leads to an underestimation of mission risk. The proposed framework successfully manages this coupled uncertainty, providing a robust and high-fidelity solution for generating practical interplanetary guidance and control algorithms.
Figure 15: 3D case: time history of spacecraft mass and its standard deviation with $\sigma_{m_{f}}\leq 40$ kg.
Figure 16: Comparison of nominal control thrust magnitudes. The controller accounting for constraint with $\sigma_{m_{f}}\leq 40$ kg (blue) requires higher thrust levels.
@@ V Conclusions
This work presents a sequential convex programming framework for solving the chance-constraint, covariance-steering problem for designing robust low-thrust interplanetary trajectories. The method incorporates a piecewise affine state-feedback policy and explicitly models spacecraft mass dynamics within the stochastic formulation. By coupling the control input and disturbance intensity through the time-varying spacecraft mass, the framework captures the evolution of uncertainty more realistically than conventional approaches. It is also possible to enforce a desired variance on the spacecraft mass at the end of the trajectory. The advantage of this model is that we can consider realistic engine parameters (such as thrust magnitude and specific impulse) and enforce constraints directly on the maximum thrust that is to be produced by the propulsion system.
Numerical results successfully demonstrate the regulation of both the mean trajectory and the state covariance in two- and three-dimensional scenarios, producing closed-loop solutions that satisfy probabilistic terminal constraints and remain consistent with Monte Carlo dispersion analysis. In particular, the inclusion of mass as a stochastic state reveals critical differences in both control effort and uncertainty evolution that are overlooked when mass is treated deterministically. Future studies will apply the computationally efficient proposed method to more challenging atmospheric [37] and cislunar [29, 38] trajectory optimization problems.